log[4]x + log[4](x+1)=log[4]6

Solve the logarithmic equation.


(To ensure my problem makes sense, the 4's are supposed to be "subscripts" of the log, I just didn't know how to type it other than this. Thank you for your help! Please respond as quickly as possible.

Answer by nerdybill(7003)

apply log rules:
log[4]x + log[4](x+1)=log[4]6
log[4](x(x+1))=log[4]6
x(x+1)=6
x^2+x = 6
x^2+x-6 =0
factoring the left:
(x+3)(x-2) = 0
x = {-3, 2}
we can throw out the -3 (extraneous solution) leaving:
x = 2